OpenStudy (anonymous):

The squares of 3 consecutive terms of an arithmetic progression are a−336, a and a+624. What is the value of a?

OpenStudy (anonymous):

Is 'a' 400?

OpenStudy (anonymous):

(middle term-d)2=a-336 middle term squared =a (middle term +common difference )squared =a+624 solve for middle term?

OpenStudy (anonymous):

yes !